# DIO: Refining Mutual Information and Causal Chain to Enhance Machine Abstract Reasoning Ability

### v1

@@ III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts DIEGO method, it demonstrates the capability to solve generative RPM problems.
The images in the RPM instances of this paper have adopted the annotation approach as illustrated in Figure 2, which enables clearer explanation of our model and method.
@@
Spcifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
@@ IV DIO: A Baseline Model
The solving process of RPM problems is as follows: The solving model needs to discern the progressive pattern of abstract image attributes from the context matrices $\{x_{i}|i\in[1,8]\}$ in an RPM instance. Subsequently, it should select the ideal option $x_{\alpha}$ from the eight options $\{x_{i}|i\in[9,16]\}$ in the option pool that will not disrupt this progressive pattern [35]. This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
This paper proposes that if the structural design of the RPM solving model is guided by this causal chain, the constructed model will demonstrate high performance. This prompted us to design the DIO model. In accordance with the causal chains in RPM problems, the DIO model is divided into four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
@@
RPM problems place significant emphasis on evaluating the abstract reasoning capabilities of deep models. Consequently, when extracting features from RPM images, we opt for a Visual Transformer (ViT) [21] that prioritizes image context. Figure 3 illustrates the process by which ViT extracts tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
@@ IV-C The Progressive Pattern Induction Module
The deliberate design of RPM problems necessitates that participants must simultaneously observe two rows and two columns to identify the progressive patterns within the matrix. This represents a critical link in the causal chain of RPM problems. Consequently, the progressive pattern induction module of DIO adopts the structure illustrated in Figure 4.
@@
Figure 4 indicates that after incorporating the features of option $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO will summarize the progressive patterns among the features that share the same token index $j$. DIO first utilizes an MLP (Multi-Layer Perceptron) to extract row and column features from the matrix. Subsequently, it randomly selects two row features from the three available row features and two column features from the three column features, combining these four features for further processing. After binding four trainable vectors of the same dimension to the selected four features, DIO employs a Transformer Encoder to process them into eight logical outputs. Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{jl}|l\in[1,4]\}$.
$\{P_{jl}|l\in[1,4]\}$ is regarded as the partial representation of progressive patterns within the matrix. If all the $C_{3}^{2}\cdot C_{3}^{2}=9$ possible ways of selecting row and column features are traversed (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), then a comprehensive representation of progressive patterns, $\{P_{jkl}|k\in[1,9],l\in[1,4]\}$, can be obtained. This module can be formally expressed as:
@@ IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations that share the same $j$-index and $l$-index. Specifically, DIO applies a Transformer Encoder to perform dimension-invariant processing on $\{P_{\alpha jkl}|k\in[1,9]\}$, resulting in $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Subsequently, each vector in $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ is mapped to a scalar via an MLP, serving as a logical score $s_{\alpha jkl}$. This process can be formally expressed as:
@@
The DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by summing $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$ according to the formula:
@@
Use the softmax function to map $\{S_{i}|i\in[8,16]\}$ to a probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$:
@@
We introduce a constraint mechanism that correlates the DIO-modeled probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$. This constraint enables joint optimization of all trainable parameters across DIO’s modules:
${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (8)
Where $\mathbb{E}(\cdot)$ denotes the expectation operation.
@@ V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
This paper proposes that architecture optimization based on causal chains can effectively enhance the performance of deep learning networks in solving RPM problems. However, through further research, we have found that this methodology still cannot fundamentally improve the logical reasoning capabilities of deep learning networks.
@@ V-A The learning objective of the DIO
By abstracting away the specific architecture of DIO (i.e., merging equations (1)-(6)), we obtain a unified formulation that highlights its core inductive mechanism:
@@
Leveraging the probabilistic framework in Equation (7), DIO can be interpreted as a tunable probabilistic model where $\theta$ represents learnable parameters:
@@
Notably, the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ can be expressed as:
@@
Where $K$ denotes the number of incorrect options. If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ , then substituting the $\hat{P}$ defined in Equation (10) into the Equation (12) yields a mutual information lower bound expressed by DIO:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right]$ (12)
When $K\geq 1$, that is, when there exists at least one incorrect option, the lower bound aligns with DIO’s optimization objective, as expressed by the following equation:
@@
Equation (13) reveals that DIO’s learning objective (i.e. maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. Furthermore, based on the properties of logarithmic functions, the following inequality holds:
${\ell}_{\text{DIO}}$ $<\mathbb{E}\left[\log\frac{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8})}\right]$
$=-1\cdot\log\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ (14)
The right-hand side of the Inequality (V-A) embodies the classical formulation for optimizing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the cross-entropy (CE) loss function, which is the typical paradigm adopted by most previous end-to-end RPM solving models. From this, it can be inferred that for those end-to-end RPM solving models that have also focused on constructing the probabilistic model $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ in the past, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (V-A) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
@@
This paper argues that employing DIO to derive a tighter variational lower bound and directing its optimization objective toward the maximization of this new lower bound can more effectively maximize mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, thereby further DIO’s performance.
By analyzing Formula (12), we deduce that increasing the number of incorrect options $K$ encountered by DIO can guide its learning objective toward a tighter variational lower bound. However, RPM problem instances only offer a limited number (seven) of incorrect options. To this end, this paper attempts to introduce additional hypothetical incorrect options for DIO. Specifically, we introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO. Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
@@
However, In Formula (16), it is not difficult to observe that if these hypothetical options cannot change DIO’s decision-making then this new lower bound, when serving as an optimization objective, will not undergo any fundamental improvement. In other words, when the probabilities assigned by the DIO model to the hypothetical options $\{x_{i}\mid i\in[17,\beta]\}$ approach zero, which means that these options do not alter the probability distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$ that DIO assigns to the original options $\{x_{i}|i\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent given that both function as loss functions for neural networks. This implies that unconstructive options contribute nothing to tightening the lower bound.
Therefore, this paper proposes introducing extra constructive hypothetical options into DIO—options that reinforce its endorsement of the correct options while simultaneously strengthening its repulsion of the incorrect ones—so as to steer DIO’s learning objective toward a tighter variational lower bound and ultimately further enhance its performance.
@@
The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options. Relying on manual design to create constructive options goes against the original intent of RPM problems. Therefore, this paper proposes training a set of constructive options instead. This methodology is referred to as Brando (Bounded Random Options Mapping Method) in this paper.
This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options for DIO, rather than their specific configurations.
Since these options are expected to improve $S_{\alpha}$ while reducing $\{S_{\hat{\alpha}}|{\hat{\alpha}}\in[9,16],{\hat{\alpha}}\neq\alpha\}$, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options and further optimize DIO:
@@
In these noval loss function:
@@
In loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options. When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.

### v3

@@ III Methodology
This paper first proposes that to enhance the ability of deep learning models to solve RPM, which are causal reasoning problems, when designing the model’s structure and dividing its modules, one should fully refer to the causal chains within the problems. Based on this methodology, this paper designs a new model for solving RPM problems, named DIO. However, this paper finds that DIO is in satin without substance. Its learning approach is essentially the same as that of the previous RPM solving model, whose purpose is to maximize the mutual information between the context matrix and the correct option in RPM instance. This learning approach leads to a situation where, no matter how well its structure aligns with the causal chains in the problems, its learning and optimization will not progress towards the human predefined reasoning framework. Therefore, this paper proposes three methods, namely Brando, WORLD, and DIEGO, to refine DIO’s learning objectives. As a bonus, when DIO adopts the WORLD method, it demonstrates the capability to solve generative RPM problems.
The RPM instances in this paper use the annotation approach illustrated in Figure 2, which allows for a clearer explanation of our model and methods.
@@
Specifically, we denote the 16 RPM images as $\{x_{i}|i\in[1,16]\}$.
@@ IV DIO: A Baseline Model
The process of solving RPM problems proceeds as follows: the model must first identify the progressive patterns governing abstract image attributes across the context matrices $\{x_{i}|i\in[1,8]\}$, and then select the correct option $x_{\alpha}$, which comprehensively continues these patterns [35], from the option pool $\{x_{i}|i\in[9,16]\}$. This implies that there exists a causal chain in RPM problems: RPM images $\rightarrow$ abstract attributes $\rightarrow$ progressive patterns of attributes $\rightarrow$ consistency of progressive patterns $\rightarrow$ correctness of options.
This paper proposes that designing the structure of an RPM solving model in accordance with this causal chain will yield high performance, which prompted us to design the DIO model. Accordingly, the DIO comprises four modules: the image feature extraction module, the progressive pattern induction module, the progressive pattern consistency evaluation module, and the option evaluation module.
@@
RPM problems place significant emphasis on evaluating the abstract reasoning capabilities of deep models. Consequently, when extracting features from RPM images, we opt for a Visual Transformer (ViT) [21] that prioritizes image context. Figure 3 illustrates the process by which ViT extracts the tokenized image features $\{z_{ij}|j\in[1,N]\}$ from the input image $x_{i}$, where $j$ denotes the token index of the image features and $N$ denotes the total patch count.
@@ IV-C The Progressive Pattern Induction Module
The deliberate design of RPM problems necessitates that participants must simultaneously observe two rows and two columns to identify the progressive patterns within the matrix. This represents a critical segment in the causal chain of RPM problems. Consequently, the progressive pattern induction module of DIO adopts the structure illustrated in Figure 4.
@@
Figure 4 indicates that after appending the option features $\{z_{\alpha j}|j\in[1,N]\}$ into the context matrix $\{z_{ij}|i\in[1,8],j\in[1,N]\}$, DIO summarizes the progressive patterns among the features that share the same token index $j$. First, DIO utilizes a Multi-Layer Perceptron (MLP) to extract row and column features from the matrix. Subsequently, it randomly selects two row features from the three available row features and two column features from the three column features, combining these four features for further processing. After binding four trainable vectors of the same dimension to the selected four features, DIO employs a Transformer Encoder to process them into eight logical outputs. Among the output vector groups, the four vectors corresponding to the trainable vectors are denoted as $\{P_{\alpha jl}|l\in[1,4]\}$.
$\{P_{\alpha jl}|l\in[1,4]\}$ are regarded as the partial representations of progressive patterns within the matrix $\{z_{ij},z_{\alpha j}|i\in[1,8]\}$. If all $C_{3}^{2}\cdot C_{3}^{2}=9$ possible row–column selections are considered (where $C_{n}^{r}$ denotes the number of combinations of $n$ items taken $r$ at a time), we obtain comprehensive representations $\{P_{\alpha jkl}|k\in[1,9],l\in[1,4]\}$ of these patterns. This module can be formally expressed as:
@@ IV-D The Progressive Pattern Consistency Evaluation Module
As its name suggests, DIO’s progressive pattern consistency evaluation module is responsible for using a Transformer Encoder to compute the consistency among progressive pattern representations $P_{\alpha jkl}$ that share the same $j$-index and $l$-index. Specifically, the Transformer Encoder performs a dimension-preserving mapping on $\{P_{\alpha jkl}|k\in[1,9]\}$ to obtain $\{P^{\prime}_{\alpha jkl}|k\in[1,9]\}$. Repeating this for all $j\in[1,N]$ and $j\in[1,4]$ yields the complete representation set $\{P^{\prime}_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$. Finally, each vector in this full set is mapped to a scalar via an MLP, producing the consistency score $s_{\alpha jkl}$. This process can be formally expressed as:
@@
DIO’s option evaluation module computes the score $S_{\alpha}$ that reflects the correctness of option $x_{\alpha}$ by averaging the consistency scores $\{s_{\alpha jkl}|j\in[1,N],k\in[1,9],l\in[1,4]\}$:
@@
The softmax function then maps $\{S_{i}|i\in[8,16]\}$ to a probability distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ parameterized by DIO:
@@
We introduce a loss function ${\ell}_{\text{DIO}}$ to refine the DIO-parameterized distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$, thereby enabling end-to-end training of all parameters across DIO’s modules:
${\ell}_{\text{DIO}}=-\mathbb{E}\left[\log\frac{\hat{P}(x=x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)}{\sum_{\hat{\alpha}=9,\hat{\alpha}\neq\alpha}^{16}\hat{P}(x=x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)}\right]$ (8)
where $\mathbb{E}[\cdot]$ denotes the expectation over the data distribution.
@@ V Deviation and Potential Correction Methods for Learning Objectives of End-to-End RPM Solving Model
This paper proposes that causal-chain-based architecture optimization can effectively enhance the performance of deep learning networks in solving RPM problems. However, further investigation reveals that this methodology still fails to fundamentally improve the logical reasoning capabilities of deep learning networks.
@@ V-A The learning objective of the DIO
By abstracting the specific architecture of DIO (i.e., by merging equations (1)–(6)), we obtain a unified formulation that highlights its core inductive mechanism:
@@
By substituting Equation (7), DIO can instantiate the parameterized probability distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$:
@@
Notably, $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, which denotes the mutual information between the context matrix $\{x_{i}|i\in[1,8]\}$ and the correct option $x_{\alpha}$, admits the following variational lower bound:
@@
where $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$ represents the true probability that $x_{\alpha}$ can serve as the correct option for the given context matrix $\{x_{i}|i\in[1,8]\}$ and $K$ denotes the number of all potential incorrect options. If we consider $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ as a sufficient approximation of the true conditional distribution $P(x_{\alpha}|\{x_{i}\}_{i=1}^{8})$, then substituting Equation (10) into Equation (11) yields a mutual information lower bound parameterized by DIO:
$I(x_{\alpha};\{x_{i}\}_{i=1}^{8})\geq\mathbb{E}\left[\log\frac{e^{\text{DIO}(\{x_{i},x_{\alpha}|i\in[1,8]\})}}{\frac{1}{K}\sum_{{\hat{\alpha}}\neq\alpha}e^{\text{DIO}(\{x_{i},x_{\hat{\alpha}}|i\in[1,8]\})}}\right],$ (12)
$\quad\quad\quad\quad\quad\quad\quad\text{where}\quad K=7$
It can be observed that the parameterized lower bound within Equation (12) aligns with DIO’s optimization objective, as expressed by the following equation:
@@
Equation (13) reveals that DIO’s learning objective (i.e. maximizing $\ell_{\text{DIO}}$) is neither the human-like reasoning logic embedded in RPM problems nor even the suboptimal approach of directly maximizing mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, but rather optimizing the variational lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Furthermore, based on the properties of logarithmic functions, the following inequality holds:
$-{\ell}_{\text{DIO}}>\mathbb{E}\left[\log\frac{\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8})}{\sum_{\hat{\alpha}=9}^{16}\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8})}\right]$ (14)
The right-hand side of Inequality (14) is precisely the classical formulation for optimizing the probabilistic distribution $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ using the Cross-Entropy (CE) loss function. From this, it can be inferred that for those past end-to-end RPM solving models that parameterize $\hat{P}(x|\{x_{i}\}_{i=1}^{8},\theta)$ and employ the CE loss function, their learning objectives essentially amount to optimizing a looser lower bound on the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$.
Equation (13) and Inequality (14) jointly uncover a pivotal fact: whether it’s the DIO model or previous end-to-end RPM solving models, their learning objectives are essentially a lower bound of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$. In this case, the looseness of the lower bound undoubtedly has a direct and significant impact on the performance of these models.
@@
This paper contends that by utilizing DIO to parameterize a tighter variational lower bound and adjusting its optimization objective to maximize this new lower bound, we can achieve a more effective maximization of the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$, ultimately enhancing DIO’s performance.
By analyzing Formula (12), we deduce that increasing the number $K$ of incorrect options encountered by DIO can guide its learning objective toward a tighter variational lower bound. However, RPM problem instances only offer a limited number (seven) of incorrect options. To this end, this paper proposes to introduce additional hypothetical incorrect options for DIO. Specifically, we attempt to introduce a set of hypothetical options $\{x_{i}|i\in[17,\beta]\}$ into the training pipeline of DIO. Consequently, the conditional distribution $\hat{P}(x_{\alpha}|\{x_{i}\}_{i=1}^{8},\theta)$ is reparameterized as:
@@
However, as shown in Formula (16), if these introduced hypothetical options fail to influence DIO’s decisions, this new lower bound, when used as an optimization objective, will not bring DIO any significant improvement. Specifically, when the probabilities that DIO assigns to the hypothetical options $\{x_{i}|i\in[17,\beta]\}$ approach zero, whereby these options cease to perturb the original distribution $\{\hat{P}(x_{\hat{\alpha}}|\{x_{i}\}_{i=1}^{8},\theta)|\hat{\alpha}\in[9,16]\}$, the variational lower bounds derived from Equation (16) and Equation (12) become equivalent when employed as loss functions. This demonstrates that unconstructive options, which are readily assigned low weights and low probabilities by DIO, contribute nothing to tightening the lower bound.
Therefore, this paper proposes introducing extra constructive hypothetical options into DIO, options which can reinforce acceptance of the correct options and simultaneously strengthen repulsion of the incorrect ones, so as to steer its learning objective toward a tighter variational lower bound and, in turn, ultimately further enhance its performance.
@@
The RPM problem expects the training process of the solving model to involve only a limited number of supervised signals, such as the index of correct options. Relying on manual design to create constructive options goes against this original intent of RPM problems. Therefore, this paper proposes training a set of constructive options instead. This methodology is referred to as Brando (Bounded Random Options Mapping Method).
This paper establishes a set of trainable hypothetical options $\{x_{i}|i\in[17,\beta]\}$ for the context matrix $\{x_{i}|i\in[1,8]\}$. This paper is not concerned with the specific attributes of these options; rather, it focuses solely on how much these options can alter DIO’s evaluation of the original options $\{x_{i}|i\in[9,16]\}$. In other words, we are solely concerned with the constructiveness of these hypothetical options with respect to the context matrix $\{x_{i}|i\in[1,8]\}$ in DIO, rather than their specific configurations. To this end, this paper designs a new loss function $\ell_{\text{Brando}}$ to train these hypothetical options for their constructiveness:
@@
In this novel loss function:
@@
In the loss function $\ell_{\text{Brando}}$, the term $\ell_{\text{up}}$ serves the purpose of tightening the variational lower bound, while $\ell_{\text{down}}$ acts as a safeguard to prevent hypothetical options from degenerating into unconstructive options. When optimizing the DIO using $\ell_{\text{Brando}}$ as the loss function, it can induce bounded hypothetical options while leveraging a tighter variational lower bound to enable DIO to maximize the mutual information $I(x_{\alpha};\{x_{i}\}_{i=1}^{8})$ more effectively.